# Building a Data Dashboard for Magic: The Gathering: Initial Design Considerations Preprint — arXiv

### v1

@@ 4.2 Winnow
After identifying the problem domain, we continued the design study with the winnow stage. During this phase, we engaged in initial discussions to identify potential collaborators within the MTG community who had a strong alignment between their domain needs and our research objectives. The Commander format helped raise an ecosystem of deck-building platforms, online discussions, and data-driven tools, making it a particularly rich case study for research in visualization and analytics. Among these tools, Gauntlet (www.gauntletapp.com), Playgroup.gg (www.playgroup.gg), Dragon Counter (www.dragoncounter.com), LifeLinked (www.sites.google.com/view/lifelinked/home), MTG Life Counter: Lotus (www.lifecounter.app) and Mythic Track (www.mythictrack.com) stand as the preferred tools to track track life, wins, or more general gameplay history. We analyzed these tools to understand the main features, data availability, and the potential to host a dashboard for Commander data. Our analysis shows that most tools do a satisfactory job of tracking generic gameplay data (e.g., damage, game record-keeping), but they either track life totals or game records. Among them, Playgroup.gg stands out as the exception, being able to track life and other gameplay data seemingly. Playgroup.gg also offers a rich history of game statistics and static dashboards with data visualizations, allowing players to explore their data and compare it against other decks, players, and playgroups (Figure 4). Therefore, we contacted the founders and developers of Playgroup.gg, Maran and Johannes, to establish a collaboration.
@@ 4.3 Cast
The subsequent cast stage involved formalizing this collaboration, confirming mutual expectations, and committing resources to jointly develop the visualization dashboard. We identified the front-line analysts as every Playgroup.gg user since they interact directly with gameplay data and would use the resulting dashboard. Additionally, the Playgroup.gg developers are the gatekeepers, who hold authority over data access and project approval. Finally, we, the authors, act as HCI researchers and translators since we need to (i) abstract the domain problems into a more generic form considering larger-context domain goals, (ii) conduct a user task analysis, and (iii) design and evaluate data visualizations to satisfy user needs related to the domain problems. Establishing these roles early ensured that we could proceed efficiently and conduct a successful iterative design of the Commander dashboard.
@@ 5 Core Phase
This section presents the stages of discover and design.
@@ 5.1.1 User-Task Analysis
We conducted a user-task analysis using a structured questionnaire designed to capture both the strengths of Playgroup.gg’s current practices and the challenges users encountered when exploring its existing static dashboards. The questionnaire was organized into four main sections to assess users’ interests and needs concerning: (i) gameplay data, (ii) relationships between different attributes, (iii) visual analytical tasks, and (iv) data comparison contexts (e.g., comparisons between a player’s decks, within a specific playgroup, or across all users). Each item was rated on a 5-point Likert scale, ranging from very uninteresting (-2) to very interesting (2). We also collected demographic data to understand our sample and assess its diversity.
@@ 5.1.2 Results
The questionnaire was released on Reddit and Discord servers focused on MTG as well as on Playgroup.gg’s webpage. A total of 68 participants filled in the questionnaire, with 58 (85.29%) currently being Playgroup.gg users. Players showed the highest interest in commonly used performance statistics, with win rate (M = 1.62, SD = 0.69) and wins versus losses (M = 1.57, SD = 0.70) rated as most important, followed by deck and player rankings (M = 1.40, SD = 0.88). Metrics related to gameplay events, such as damage and healing (M = 1.04, SD = 1.16), number of rounds (M = 0.99, SD = 0.97), and starting player position (M = 0.87, SD = 1.13), were moderately interesting, whereas less common statistics like estimated deck power level (M = 0.82, SD = 1.23), kill frequency (M = 0.76, SD = 0.92), and ELO (M = 0.71, SD = 1.21) received lower interest. ELO is a method for calculating the relative skill levels of players in zero-sum games. Players expressed little to negative interest in peripheral or largely chance-driven metrics such as the number of priority changes (M = -0.13, SD = 1.21), number of dice rolls (M = -0.38, SD = 1.21), and pause duration (M = -0.63, SD = 1.08). Overall, these results suggest that users prioritize metrics directly related to game outcomes and player/deck performance over auxiliary or stochastic elements of gameplay.
@@
Players showed the greatest interest in dashboards that provide a Playgroup view, focusing on members of a specific playgroup (M = 1.32, SD = 0.95), followed by a User view, which isolates data from the individual player’s decks (M = 1.07, SD = 1.20). The Global view, aggregating data across all playgroups, was considerably less interesting to users (M = 0.21, SD = 1.31), suggesting that players value contextually relevant comparisons over broad, population-level analyses. These values are in line with the view of Commander as a casual and social format. Regarding insights, users prioritized analyses that reveal strategically actionable patterns, with the highest interest in exploring the relationship between games played, games won, and the number of kills per commander (M = 1.22, SD = 1.08). Other moderately interesting insights included the popularity of color identities and commanders (M = 0.97, SD = 1.09), while metrics such as time spent per turn, damage per turn by commander, and damage type distributions per commander received lower interest (M = 0.46–0.71).
Afterwards, we analyzed the qualitative data, where players suggested a wide range of additional metrics, analyses, and insights that could enhance the visualization of Commander gameplay. Several ideas focused on deck- or player-level performance, including metrics like archenemy tracking to measure which players or decks are perceived as threats, average turn to cast a commander, or commander casts per game. Other gameplay aspects included identifying trends in personal skill and deck effectiveness. Regarding analysis types, players expressed interest in metrics that could reveal actionable patterns such as the ability to compare performance across different playgroup sizes or against players with differing skill levels. Overall, these results indicate that players value strategically actionable, turn-sensitive, and contextually relevant metrics, with a preference for views and insights that emphasize comparisons within their own playgroup or deck context. They also seek the flexibility to track long-term trends, deck evolution, and analyses that connect directly to game outcomes and strategic performance.
@@ 5.2 Design
In the design stage, we translated the insights gathered during the discover phase into concrete visualization solutions for Playgroup.gg. This process involved generating and validating data abstractions, visual encodings, and interaction mechanisms that could effectively communicate gameplay and strategic information to players. Before narrowing the proposal space in consultation with the Playgroup.gg founders, we iteratively explored multiple design alternatives, including different chart types, filtering options, and comparison contexts (e.g., individual decks, playgroup-wide, or global views). We ran two rapid prototyping workshops to create sketches and mockups of an interactive dashboard to enable quick feedback loops, allowing us to refine visual encodings and interaction features based on domain expert input.
@@ 5.2.1 The Dashboard
Figure 5 presents the dashboard mockup the authors proposed to the Playgroup.gg team. The dashboard provides a comprehensive overview of player and deck performance in the Commander format, integrating multiple coordinated visualizations to support analysis at both granular and aggregate levels. The top section features a parallel coordinates plot summarizing key metrics such as number of wins and games played, damage, healing, ELO, and estimated deck power, enabling users to quickly compare multiple commanders across these dimensions. To the right, bar charts illustrate the frequency of specific game end states, facilitating the identification of deck strategies. Below, a scatterplot maps games played against games won, with shading encoding average kills, providing a clear view of performance trends, while an adjacent line chart tracks damage dealt per turn by commanders, supporting temporal analysis. The lower-left heatmap shows seating positions across multiple metrics, allowing rapid assessment of relative performance by play order. Finally, an icicle plot at the bottom-right encodes the number of decks per color identities and number of colors with an extra brightness encoding for player “saltiness”, highlighting patterns in player experiences and specific color combinations and commanders. Finally, the dashboard includes panels for active filters and customization instructions, ensuring that users can tailor views to specific analytical questions and comparative contexts.
Figure 5: Mockup of the proposed dashboard for Commander data analysis for Playgroup.gg.
@@ 5.2.2 Evaluation
We then met with the Playgroup.gg team to gather their feedback on the proposed dashboard. While they praised the overall design, they noted one weakness: using only a black stroke for the marks in the line chart made it difficult to track a single line across the x-axis. After consideration, we decided to retain the black stroke, as the goal of the visualization was not to individually identify commanders since this functionality was already supported through filtering in the right-hand panel. Instead, we want to analyze the overall evolution of damage per turn. Following this, we collaborated with the Playgroup.gg team to develop a questionnaire designed to assess whether users could understand and extract insights from the charts. The questionnaire contained one section per chart, with each item structured as a multiple-choice question with only one correct answer. Each item required participants to interpret a chart and respond accordingly. To reduce random guessing, participants were explicitly instructed not to guess and were given the option to select I don’t know if undecided. We framed the questionnaire as a quiz, awarding one point per correct answer, to foster engagement. Finally, participants were invited to provide open-ended feedback as well as demographic information to contextualize the sample.
@@ 5.2.3 Results
Many participants requested interactive features, such as hovering or clicking to reveal precise values, filtering by attributes (e.g., power level, game length), and brushing/selection across linked charts. Customization was also a recurring theme. Users wanted the ability to “remove/add charts at your will,” “change each graph to show different states,” or select their preferred chart type for a given metric. This reflects a desire for dashboards that adapt to individual preferences and reduce information overload. Several participants also noted that the dashboard felt “too information dense” and “intimidating” when all charts were displayed at once. They suggested breaking the dashboard into sections, displaying charts one at a time, or providing quick-glance summaries (e.g., “top 3 winrates” or “saltiest commanders”) before drilling into detail. Additionally, participants mentioned the need for clearer separation between charts (“hard edges to separate each chart”) and consistency in elements like legend placement.
@@
5.3 Reflect
@@
The results of this study allow us to confirm, refine, and propose design guidelines for dashboards supporting gameplay analysis in digital platforms such as Playgroup.gg. First, our findings confirm existing guidelines emphasizing the importance of contextually relevant, outcome-driven metrics: players consistently valued statistics like win rate and wins versus losses, while peripheral or stochastic measures such as dice rolls or pause duration were rejected as uninformative. Second, we refine guidelines for visualization choice by showing that although canonical charts such as heatmaps, parallel coordinates, and line charts supported high comprehension (M $\geq$ 78.9%), scatterplots and icicle plots posed substantial interpretation challenges, particularly under non-ideal presentation conditions such as black-and-white rendering. These results suggest the need to refine design guidance around complex or unconventional charts, recommending their use only when paired with interactivity (e.g., hover-to-reveal, filtering, brushing) or simplified encodings. Third, our findings reject the assumption that global or population-level dashboards are inherently useful: participants strongly preferred playgroup- and user-level views, suggesting that more aggregated, population-level data provides broader insight and is therefore more valuable. Finally, we propose new guidelines emphasizing user customization and progressive disclosure. Participants frequently requested the ability to tailor which charts were shown, select preferred encodings, and control dashboard density (e.g., quick summaries before detail), underscoring that adaptability is as critical as accuracy in supporting diverse player needs. Together, these reflections extend design study knowledge by grounding visualization guidelines in the practices and expectations of a specific yet illustrative community, while pointing to generalizable principles for balancing interpretability, customization, and contextual relevance in dashboard design.
@@
Future work should review the proposed design in light of our results and continue with a prototype implementation and deployment in the Playgroup.gg website. Additionally, researchers may further investigate how interactive features can mitigate the challenges observed with complex visualizations such as scatterplots and icicle plots, particularly in mobile or constrained environments where overlapping elements are harder to trace. Further studies could explore adaptive dashboards that personalize chart selection and level of detail to different user profiles, such as novice players seeking simple summaries versus expert players interested in advanced cross-metric analyses. Incorporating player-centered enhancements such as commander images, mana symbols, or icons may improve recognition and engagement, and systematic testing of accessibility features (e.g., colorblind palettes, light/dark modes) could yield generalizable guidelines. Importantly, future research should broaden the participant base beyond Playgroup.gg users. Since our sample primarily consisted of individuals already interested in statistics and data tracking, their preferences may not fully represent the general Commander audience. Comparing results across both data-inclined and casual player populations would clarify whether the observed prioritization of strategic, outcome-driven metrics reflects a universal trend or a community-specific bias. Finally, longitudinal research is needed to understand how player preferences and comprehension evolve over time as familiarity with the dashboard grows, helping to refine guidelines around learning curves and long-term adoption.

### v2

@@ 4.2 Winnow
After identifying the problem domain, we continued the design study with the winnow stage. During this phase, we engaged in initial discussions to identify potential collaborators within the MTG community who had a strong alignment between their domain needs and our research objectives. The Commander format helped raise an ecosystem of deck-building platforms, online discussions, and data-driven tools, making it a particularly rich case study for research in visualization and analytics. Among these tools, Gauntlet [gauntletappGauntlet], Playgroup.gg [playgroug], Dragon Counter [dragoncounterDragonCounter], LifeLinked [lifecounterLifeCounter], MTG Life Counter: Lotus [lifecounterLifeCounter] and Mythic Track [mythictrackTrackYour] stand as the preferred tools to track track life, wins, or more general gameplay history. We analyzed these tools to understand the main features, data availability, and the potential to host a dashboard for Commander data. Our analysis shows that most tools do a satisfactory job of tracking generic gameplay data (e.g., damage, game record-keeping), but they either track life totals or game records. Among them, Playgroup.gg stands out as the exception, being able to track life and other gameplay data seemingly. Playgroup.gg also offers a rich history of game statistics and static dashboards with data visualizations, allowing players to explore their data and compare it against other decks, players, and playgroups (subsection 4.3). Therefore, we contacted the founders and developers of Playgroup.gg, Maran and Johannes, to establish a collaboration.
@@ 4.3 Cast
The subsequent cast stage involved formalizing this collaboration, confirming mutual expectations, and committing resources to jointly develop the visualization dashboard. We identified the front-line analysts as every Playgroup.gg user since they interact directly with gameplay data and would use the resulting dashboard. Additionally, the Playgroup.gg developers are the gatekeepers, who hold authority over data access and project approval. Finally, we, the authors, act as HCI researchers and translators since we need to (i) abstract the domain problems into a more generic form considering larger-context domain goals, (ii) conduct a user task analysis, and (iii) design and evaluate data visualizations to satisfy user needs related to the domain problems. Establishing these roles early ensured that we could proceed efficiently and conduct a successful iterative design of the dashboard.
figure*[!t] [Deck dashboard.][Uncaptioned image] [Player dashboard.][Uncaptioned image] [Playgroup dashboard.][Uncaptioned image] Screenshots of the dashboards developed by Playgroup.gg.
@@ 5 Core Phase
The second phase of the design study methodology comprises four stages: discover, design, implement, and deploy. In this study, we completed only the first two stages; the remaining stages are reserved for future work.
@@ 5.1.1 User-Task Analysis
We conducted a user-task analysis using a structured questionnaire designed to capture both the strengths of Playgroup.gg’s current practices and the challenges users encountered when exploring its existing static dashboards. Although Sedlmair et al. mention interviews as a viable method [sedlmair2012design] and that interviews generate better insights than surveys [Zunaed2021], we decided to retain the questionnaires. Questionnaires allow large amounts of information to be summarized efficiently [Reid1988Questionnaires], reach much larger populations [Zunaed2021], and generate data that can be easily quantified and statistically analyzed [Reid1988Questionnaires]. The questionnaire was organized into four main sections to assess users’ interests and needs concerning: (i) gameplay data, (ii) relationships between different attributes, (iii) visual analytical tasks, and (iv) data comparison contexts (e.g., comparisons between a player’s decks, within a specific playgroup, or across all users). Each item was rated on a 5-point Likert scale, ranging from very uninteresting (-2) to very interesting (2). We also collected demographic data to understand our sample and assess its inner diversity.
@@ 5.1.2 Results
The questionnaire was released on Reddit and Discord servers focused on MTG as well as on Playgroup.gg’s webpage. We wanted to ensure that most of our participants were familiar with MTG to avoid any bias due to domain misunderstandings. Our results were analyzed using only descriptive statistics, as they are useful for uncovering patterns within the data and providing simple summaries about the sample and the measures [Nick2007, Gupta2019]. A total of 68 participants filled in the questionnaire, with 58 (85.29%) currently being Playgroup.gg users. Players showed the highest interest in commonly used performance statistics, with win rate (M = 1.62, SD = 0.69) and wins versus losses (M = 1.57, SD = 0.70) rated as most important, followed by deck and player rankings (M = 1.40, SD = 0.88). Metrics related to gameplay events, such as damage and healing (M = 1.04, SD = 1.16), number of rounds (M = 0.99, SD = 0.97), and starting player position (M = 0.87, SD = 1.13), were moderately interesting, whereas less common statistics like estimated deck power level (M = 0.82, SD = 1.23), kill frequency (M = 0.76, SD = 0.92), and ELO (M = 0.71, SD = 1.21) received lower interest. ELO is a method for calculating the relative skill levels of players in zero-sum games. Players expressed little to negative interest in peripheral or largely chance-driven metrics such as the number of priority changes (M = -0.13, SD = 1.21), number of dice rolls (M = -0.38, SD = 1.21), and pause duration (M = -0.63, SD = 1.08). Overall, these results suggest that users prioritize metrics directly related to game outcomes and player/deck performance over auxiliary or stochastic elements of gameplay.
@@
Players showed the greatest interest in dashboards that provide a Playgroup view, focusing on members of a specific playgroup (M = 1.32, SD = 0.95), followed by a User view, which isolates data from the individual player’s decks (M = 1.07, SD = 1.20). The Global view, aggregating data across all playgroups, was considerably less interesting to users (M = 0.21, SD = 1.31), suggesting that players value contextually relevant comparisons over broad, population-level analyses. These values are in line with the view of Commander as a casual and social format. Regarding insights, users prioritized analyses that reveal strategically actionable patterns, with the highest interest in exploring the relationship between games played, games won, and the number of kills per commander (M = 1.22, SD = 1.08). Other moderately interesting insights included the popularity of color identities and commanders (M = 0.97, SD = 1.09), while metrics such as the time spent per turn, damage per turn by commander, and damage type distributions per commander received lower interest (M = 0.46–0.71).
Afterward, we analyzed the qualitative data, where players suggested a wide range of additional metrics, analyses, and insights that could enhance the visualization of Commander gameplay. Several ideas focused on deck- or player-level performance, including metrics like archenemy tracking to measure which players or decks are perceived as threats, average turn to cast a commander, or commander casts per game. Other gameplay aspects included identifying trends in personal skill and deck effectiveness. Regarding analysis types, players expressed interest in metrics that could reveal actionable patterns such as the ability to compare performance across different playgroup sizes or against players with differing skill levels. Overall, these results indicate that players value strategically actionable, turn-sensitive, and contextually relevant metrics, with a preference for views and insights that emphasize comparisons within their own playgroup or deck context. They also seek the flexibility to track long-term trends, deck evolution, and analyses that connect directly to game outcomes and strategic performance.
The following high-level objectives were established: comparative analysis (T1), relationship mapping (T2), trend identification (T3), and hierarchical evaluation (T4). These objectives shaped the methodology for the subsequent stage.
@@ 5.2 Design
During the design stage, we translated insights from the discovery phase into concrete visualization solutions for Playgroup.gg, beginning the development of our mock design implementation. This process involved generating and validating data abstractions, visual encodings, and interaction mechanisms that could effectively communicate gameplay and strategic information to players. Before narrowing the proposal space in consultation with the Playgroup.gg founders, we iteratively explored multiple design alternatives, including different chart types, filtering options, and comparison contexts (e.g., individual decks, playgroup-wide, or global views). We ran two rapid prototyping workshops to create sketches and mockups of an interactive dashboard to enable quick feedback loops, allowing us to refine visual encodings and interaction features based on domain expert input.
@@ 5.2.1 The Dashboard
Figure 5 presents the dashboard mockup the authors proposed to the Playgroup.gg team. The dashboard provides a comprehensive overview of player and deck performance in the Commander format, integrating multiple coordinated visualizations to support analysis at both granular and aggregate levels. The top section features a parallel coordinates plot (T1 and T2) summarizing key metrics such as number of wins and games played, damage, healing, ELO, and estimated deck power, as this type of chart enables users to quickly compare multiple commanders across several dimensions [Blaas2008]. To the right, bar charts (T1) illustrate the frequency of specific game end states, facilitating the identification of deck strategies. Below, a scatterplot (T2) maps games played against games won, with shading encoding average kills, providing a clear view of performance trends, while an adjacent line chart (T3) tracks damage dealt per turn by commanders, supporting temporal analysis. The lower-left heatmap (T1 and T2) shows seating positions across multiple metrics, allowing rapid assessment of relative performance by play order. Finally, an icicle plot (T4) at the bottom-right encodes the number of decks per color identities and number of colors with an extra brightness encoding for player “saltiness”, highlighting patterns in player experiences and specific color combinations and commanders. We chose this type of chart because it performs well compared to hierarchical alternatives such as treemaps [Macquisten2020-in]. Finally, the dashboard includes panels for active filters and customization instructions, ensuring that users can tailor views to specific analytical questions and comparative contexts. Our next step was to evaluate the dashboard. We retained it as a static, black-and-white prototype, since research indicates that prototype medium has no significant main effect on users’ ability to detect usability problems or on their perceptions of usability [Boothe2013-zx].
Figure 5: Mockup of the proposed dashboard for Commander data analysis for Playgroup.gg. The top section features a parallel coordinates plot summarizing key metrics, with bar charts to the right showing the frequency of specific game end states. Below, a scatterplot maps games played against games won, while an adjacent line chart tracks damage dealt per turn by commanders. The lower-left heatmap displays seating positions across multiple metrics. At the bottom-right, an icicle plot encodes deck counts by color identity and number of colors, with brightness indicating player ’saltiness’ (a measure of frustration or playstyle intensity). Finally, the dashboard includes panels for active filters and customization instructions.
@@ 5.2.2 Evaluation
We met with the Playgroup.gg team to gather their feedback on the proposed dashboard. While they praised the overall design, they noted one weakness: using only a black stroke for the marks in the line chart made it difficult to track a single line across the x-axis. After consideration, we decided to retain the black stroke, as the goal of the visualization was not to individually identify commanders since this functionality was already supported through filtering in the right-hand panel. Instead, we want to analyze the overall evolution of damage per turn. Following this, we collaborated with the Playgroup.gg team to develop a questionnaire designed to assess whether users could understand and extract insights from the charts. The questionnaire contained one section per chart, with each item structured as a multiple-choice question with only one correct answer. Each item required participants to interpret a chart and respond accordingly. To reduce random guessing, participants were explicitly instructed not to guess and were given the option to select I don’t know if undecided. We framed the questionnaire as a quiz, awarding one point per correct answer, to foster engagement. Finally, participants were invited to provide open-ended feedback as well as demographic information to contextualize the sample.
@@ 5.2.3 Results
Many participants requested interactive features, such as hovering or clicking to reveal precise values, filtering by attributes (e.g., power level, game length), and brushing/selection across linked charts. Customization was also a recurring theme. Users wanted the ability to “remove/add charts at your will,” “change each graph to show different states,” or select their preferred chart type for a given metric. This reflects a desire for dashboards that adapt to individual preferences and reduce information overload. Several participants also noted that the dashboard felt “too information dense” and “intimidating” when all charts were displayed at once. We believe that this happened because data visualization literacy among adults in the general population is moderate to low, with substantial variation across chart types and demographic groups [Galesic2010]. Participants thus suggested breaking the dashboard into sections, displaying charts one at a time, or providing quick-glance summaries (e.g., “top 3 winrates” or “saltiest commanders”) before drilling into detail. Additionally, participants mentioned the need for clearer separation between charts (“hard edges to separate each chart”) and consistency in elements like legend placement.
@@
6 Analysis
The final phase of the design study methodology, which represents the discussion of our work, consists of two stages: reflect and write. The write stage produced this paper as its primary output. The reflect stage encompasses all preceding stages except for implementation and deployment.
@@ 6 Analysis
The results of this study allow us to confirm, refine, and propose design guidelines for dashboards supporting gameplay analysis in digital platforms such as Playgroup.gg. First, our findings confirm existing guidelines emphasizing the importance of contextually relevant, outcome-driven metrics: players consistently valued statistics like win rate and wins versus losses, while peripheral or stochastic measures such as dice rolls or pause duration were rejected as uninformative. Second, we refine guidelines for visualization choice by showing that although canonical charts such as heatmaps, parallel coordinates, and line charts supported high comprehension (M $\geq$ 78.9%), scatterplots and icicle plots posed substantial interpretation challenges, particularly under non-ideal presentation conditions such as black-and-white rendering. These results suggest the need to refine design guidance around complex or unconventional charts, recommending their use only when paired with interactivity (e.g., hover-to-reveal, filtering, brushing) or simplified encodings. Third, our findings reject the assumption that global or population-level dashboards are inherently useful: participants strongly preferred playgroup- and user-level views, suggesting that more aggregated, population-level data provides broader insight and is therefore more valuable. Finally, we propose new guidelines emphasizing user customization and progressive disclosure. Participants frequently requested the ability to tailor which charts were shown, select preferred encodings, and control dashboard density (e.g., quick summaries before detail), underscoring that adaptability is as critical as accuracy in supporting diverse player needs. Together, these reflections extend design study knowledge by grounding visualization guidelines in the practices and expectations of a specific yet illustrative community while pointing to generalizable principles for balancing interpretability, customization, and contextual relevance in dashboard design.
In the end, we successfully applied the pre-condition and core phases of the design study methodology proposed by Sedlmair et al. [sedlmair2012design], a methodological extension of PD [frishberg2011interactive, kahng2017cti], to another rich data visualization domain: MTG. In doing so, we contributed to InfoVis by introducing a TCG to data visualization for games, a game type not previously explored with data exploration prototypes. We hope our solution may improve player behavior and experience in Commander, as similar interventions have already demonstrated in other games [Wallner2014]. Although MTG has been explored by the community [githubGitHubVincenzoGalantemagicthegathering], there was a lack of sound research methodologies and focus on gameplay, which we aimed to address. By formally documenting and explaining the stage details presented in the previous sections, we hope the InfoVis community can use our use case as an example for other domains, whether TCGs or other game types. Finally, while we carefully followed user needs to propose our dashboard, we hope our ideas spark curiosity about learning new visual techniques to represent data. This demonstrates patterns that many players remain unaware of [Goyal2013] due to low visualization literacy [Galesic2010], while ensuring Playgroup.gg creators see their collected data presented in creative and innovative ways, showcasing the power of InfoVis.
7 Future Work
@@
The proposed design should be reviewed in light of our results and continue with a prototype implementation and deployment in the Playgroup.gg website. Additionally, researchers may further investigate how interactive features can mitigate the challenges observed with complex visualizations such as scatterplots and icicle plots, particularly in mobile or constrained environments where overlapping elements are harder to trace. Further studies could explore adaptive dashboards that personalize chart selection and level of detail to different user profiles, such as novice players seeking simple summaries versus expert players interested in advanced cross-metric analyses. Incorporating player-centered enhancements such as commander images, mana symbols, or icons may improve recognition and engagement, and systematic testing of accessibility features (e.g., colorblind palettes, light/dark modes) could yield generalizable guidelines. Importantly, future research should broaden the participant base beyond Playgroup.gg users. Since our sample primarily consisted of individuals already interested in statistics and data tracking, their preferences may not fully represent the general Commander audience. Comparing results across both data-inclined and casual player populations would clarify whether the observed prioritization of strategic, outcome-driven metrics reflects a universal trend or a community-specific bias. Finally, longitudinal research is needed to understand how player preferences and comprehension evolve over time as familiarity with the dashboard grows, helping to refine guidelines around learning curves and long-term adoption.